OpenStudy (anonymous):

Find the distance between P1(4,16degrees) and P2(-2,177degrees) on the polar plane.

OpenStudy (anonymous):

\[\sqrt{(4+2)^2+(16-177)^2}\]

OpenStudy (anonymous):

\[\sqrt{(36+37249)}\]

OpenStudy (anonymous):

\[\sqrt{37285}\]

OpenStudy (anonymous):

\[5\sqrt{7457}\]

OpenStudy (anonymous):

I think that's as far as you can reduce it I think that's it's most radical form

OpenStudy (anonymous):

sorry that's the best i could do

OpenStudy (anonymous):

IDFWNOONE

OpenStudy (anonymous):

http://gamutech.com